# Valeo4Cast: A Modular Approach to End-to-End Forecasting

### v1

@@ 4 Conclusion
The modular pipeline Valeo4Cast ranks first in the AV2 E2E Forecasting challenge 2024 and outperforms other solutions by +13 mAP pts. This design allowed us to start from a state-of-the-art motion forecasting model, that we integrate into an end-to-end pipeline by finetuning the forecasting module. We include a post-processing step to account for the absence of static objects in the conventional motion forecasting pretraining but which are important in the end-to-end benchmark. In this work, we confirm the findings of [27], verifying the superiority of modular approaches in the end-to-end forecasting task, and their capacity to handle detection and tracking imperfections.
The efficient nature of the end-to-end approaches is still appealing. In future work, we are interested in investigating how to better train the end-to-end approaches in order to achieve performances on-par with Valeo4Cast. Besides, future work may also consider more challenging settings in which the map information is not provided, at any stage, and has to be inferred in an online fashion, as the ego car drives and discovers its environment.
Acknowledgment
The project was partially funded by ANR grant MultiTrans (ANR-21-CE23-0032). This research received the support of EXA4MIND project, funded by a European Union’s Horizon Europe Research and Innovation Programme, under Grant Agreement N°101092944 views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. The project also receives compute resources from the petascale system Karolina, acquired as part of the EuroHPC. We thank the authors of UniTraj [6] for the release of the forecasting framework. We also thank the organizers of the challenge.

### v2

@@ 5 Conclusion
The modular pipeline Valeo4Cast ranks first in the AV2 E2E Forecasting challenge 2024 and outperforms other solutions by 13 mAP pts. This design allowed us to start from a state-of-the-art motion forecasting model, that we integrate into an end-to-end pipeline by finetuning the forecasting module. We conduct in-depth study on how to finetune the forecaster with different matching strategies. We include a post-processing step to account for the absence of static objects in conventional motion forecasting pretraining which are important in the end-to-end benchmark. In this work, we confirm the findings of [29], verifying the superiority of modular approaches in the end-to-end forecasting task, and their ability to handle detection and tracking imperfections.
The efficient nature of the end-to-end approaches is still appealing. In future work, we are interested in investigating how to better train the end-to-end approaches to achieve performances on par with Valeo4Cast. Besides, future work may also consider more challenging settings in which the map information is not provided, at any stage, and has to be inferred in an online fashion, as the ego car drives and discovers its environment.